A park has a permiter of 140. If the length of the park is 20meters more than its width.,what is the are?

Answer by phillywily(18)

Take width as x, and since length is 20 longer than width, length should be (x+20).
Perimeter=140
140=2(x+20)+2x
140=2x+40+2x
140=4x+40
140-40=4x
100=4x
x=25
Width=25
Length
=25+20
=45
Area
=Width×Length
=25(45)
=1125m²